OpenStudy (anonymous):

A company has purchased 5 company trucks for $23,000 each. If the depreciation rate is 5.75%, what is the amount that depreciated per vehicle after 4 years? A. $4,850.98 B. $18,149.02 C. $24,254.88 D. $90,745.12

OpenStudy (anonymous):

i think its C. $24,254.88

OpenStudy (texaschic101):

1st year -- 23,000 x .0575 = 1322.50 23,000 - 1322.5 = 21,677.50 2nd year -- 21,677.50 x .0575 = 1246.45 21,677.50 - 1246.45 = 20,431.05 3rd year -- 20,431.05 x .0575 = 1174.79 20,431.05 - 1174.79 = 19,256.26 4th year -- 19,256.26 x .0575 = 1107.23 19,256.26 - 1107.23 = 18,149.03 I am off by one penny....must have made a mistake in rounding and I am too lazy to find it....regardless, the answer is B.